# RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer

### v1

@@ Abstract
Imitation Learning has become a fundamental approach in robotic manipulation. However, collecting large-scale real-world robot demonstrations is prohibitively expensive. Simulators offer a cost-effective alternative, but the sim-to-real gap make it extremely challenging to scale. Therefore, we introduce RoboTransfer, a diffusion-based video generation framework for robotic data synthesis. Unlike previous methods, RoboTransfer integrates multi-view geometry with explicit control over scene components, such as background and object attributes. By incorporating cross-view feature interactions and global depth/normal conditions, RoboTransfer ensures geometry consistency across views. This framework allows fine-grained control, including background edits and object swaps. Experiments demonstrate that RoboTransfer is capable of generating multi-view videos with enhanced geometric consistency and visual fidelity. In addition, policies trained on the data generated by RoboTransfer achieve a 33.3% relative improvement in the success rate in the Diff-Obj setting and a substantial 251% relative improvement in the more challenging Diff-All scenario. Explore more demos on our project page: https://horizonrobotics.github.io/robot_lab/robotransfer
**footnotetext: Equal contribution. footnotetext: Corresponding author.
@@ 1 Introduction
Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [67, 68]. However, collecting large-scale real-world robot demonstrations is prohibitively expensive [6, 7]. Simulated environments offer a cost-effective alternative [54, 59, 19, 44], but the scarcity of assets and sim-to-real gap[51, 43] make it extremely challenging to scale.
@@
Despite the rapid advancements in the field, two key challenges remain. First, robotic systems often rely on observations from a single viewpoint, which inherently limits their perception. To overcome this limitation, multi-view observation is a commonly used approach. However, generating multi-view consistent results remains a significant challenge for video generative models. Second, robot manipulation scenarios are typically complex and highly interactive. Controlling such scenarios precisely through textual input alone, as demonstrated in text-to-video frameworks [5, 70], presents substantial difficulties.
In this work, we introduce RoboTransfer, as is shown in Figure 1, a video generation framework designed for data synthesis in robotics. To the best of our knowledge, it is the first framework to integrate multi-view geometry while offering explicit control over background, object attributes. RoboTransfer incorporates cross-view feature interactions and global depth/normal conditions within the video diffusion architecture, ensuring 3D consistency across different camera views and over time. To enhance controllability, RoboTransfer decomposes each scene into distinct control conditions, enabling background edits and object swaps by conditioning on this decomposition.
[…]
Figure 1: Overview of RoboTransfer.
Our work makes the following contributions:
@@
    We propose RoboTransfer, a robot generation framework based on multi-view, geometry-aware video diffusion models that maintain fidelity across multiple reference conditions.
@@
    We introduce a dataset construction pipeline that decomposes real-world recording data into high-quality triplets, which incorporate global geometry and appearance control conditions.
@@
    Through carefully designed experiments and evaluations, we demonstrate that our approach generates multi-view videos with significantly improved geometric consistency and visual fidelity. Furthermore, policies trained with RoboTransfer-generated data exhibit better generalization to novel and unseen environments.
@@ 2.1 Visual Generalizable Imitation Learning
Imitation Learning (IL) has emerged as a foundational approach for visuomotor control in robotic manipulation. Deep networks trained on collected demonstrations—e.g., DP [12], OpenVLA [31], Act [67], Aloha [68], E2E-VP [36], and RT2 [7]—map raw visual inputs directly to motor commands. To bolster policy robustness and generalization, researchers have explored novel architectures [12] and vision pretraining schemes [4, 13]. However, gathering large-scale, task-specific robot data remains prohibitively expensive. Consequently, many methods leverage auxiliary sources—human demonstration videos [20], high-fidelity simulators [54, 59, 19, 44], or operational logs from other platforms [46, 30]. Each of these domains exhibits its own biases and visual characteristics, introducing distributional shifts that can degrade IL policy performance when transferred to the target environment. In this work, we offer a complementary perspective on generalizable visual imitation learning by leveraging state-of-the-art video diffusion models. By generating target-conditioned training data, our approach significantly enhances the robustness of policy and enables seamless transfer to novel environments.
2.2 Data Augmentation for Policy Learning
Data augmentation is widely used to train robust machine learning models. Prior work has applied augmentation directly to real-world robot captures to improve the resilience of visual policies against environmental variations[41, 42, 15, 23, 21]. Image-inpainting methods further enhance appearance invariance by replacing or masking out scene elements in collected frames [62, 11, 10, 40]. Likewise, neural view synthesis techniques generate novel viewpoints from existing robot data, promoting view-invariant feature learning [53, 8]. Simulators can cheaply produce vast quantities of labeled frames [54, 59, 19, 44], but discrepancies in physics modeling, rendering fidelity, and scene composition often hinder direct sim-to-real transfer. Domain randomization is used to close the sim-to-real gap, thereby enabling policies trained in simulation to generalize more effectively to real-world environments. [35, 22, 34, 51, 1, 48, 25]. In contrast to these methods, we leverage a video diffusion model to generate photorealistic image manipulations. This approach preserves both multi-view and geometry consistency, producing richly detailed samples for robot policy training.
@@ 2.3 Generation Models in Embodied AI
With the rapid advancement of generative models such as HunyuanVideo [33], Wan [56], SVD [5], Cosmos [2], CogVideoX [61], OpenSora [70], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [72]. In particular, domains like autonomous driving have seen increasing adoption of generative techniques, as demonstrated by [57, 66, 27, 58, 17, 18, 64, 45, 65]. Similarly, robotic scenarios—where real-world data collection is often prohibitively expensive and time-consuming—have greatly benefited from generative data. Models such as UniSim [60], UniPi [14], and RoboDreamer [71] synthesize future robot behaviors via text prompts, which are then translated into actionable commands using inverse dynamics models. However, purely video-based generation methods often suffer from spatial and temporal inconsistencies, leading to degraded performance in downstream action prediction and planning. To address this, TesserAct [69] proposes a multi-modal generation pipeline that simultaneously synthesizes RGB, depth, and normal videos, enabling the reconstruction of coherent 4D scenes. This framework ensures both spatial and temporal consistency in robotic environments and supports policy learning that significantly surpasses prior video-only world models. Moreover, data generalization remains a critical challenge in robot learning. Techniques like ReBot [16] and RoboEngine [63] adopt background inpainting to diversify environmental textures, while Cosmos-Transfer [2] utilizes video-to-video translation to enrich overall scene appearance. Despite their success, these approaches lack fine-grained control over foreground and background textures, often resulting in synthetic data distributions misaligned with real-world tasks, especially when training policy models. Additionally, their inability to ensure multi-view consistency limits their scalability and effectiveness in complex, robotic settings.
@@ 3.1 Video Diffusion Model Preliminaries
Diffusion-based methods for controllable video generation typically model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map, segmentation map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding). Popular approaches in this paradigm include Flow Matching [37], DDPM [26], and EDM [29]. In the case of EDM, the learning objective is formulated as:
$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$ (1)
@@
$\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
@@
$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).$ (4)
@@
Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints. It represents geometry with scaled depth and normal maps, and encodes appearance using reference backgrounds and object-specific images for detailed control over appearance.
3.2 RoboTransfer Framework
@@
The overall framework of RoboTransfer is illustrated in Figure 2. To ensure multi-view consistency during generation, we perform multi-view consistent modeling, enabling the generation process to reason jointly over information from different viewpoints. On the conditions side, RoboTransfer incorporates fine-grained control by encoding both geometric and appearance information. Specifically, we represent geometry using scaled depth maps and surface normal maps, capturing the underlying 3D structure of the scene. Meanwhile, the appearance is encoded using reference background images and object-specific images, providing detailed control over texture, color, and contextual appearance. In the following sections, we first introduce the multi-view consistent modeling, and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
@@
Multi-view Consistent Modeling. In robotic environments, it is common to deploy multiple cameras from different viewpoints to capture a scene simultaneously. To generate multi-view consistent videos for downstream tasks, RoboTransfer leverages the multi-view in-context learning capabilities [66, 28] inherently present in pretrained diffusion models. Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a VAE encoder $\mathcal{E}$:
$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),$ (5)
@@
where $[V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence. This modeling approach maintains the integrity of the pretrained video diffusion architecture, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic settings.
@@
Geometry Conditions. Recent video generation models primarily learn spatiotemporal coherence through data-driven approaches such as temporal attention [5, 70] or full attention [56, 33, 2]. While effective at capturing pixel-level dynamics, these methods lack an explicit understanding of underlying 3D geometry, limiting their applicability in robotic environments. To address this limitation, RoboTransfer explicitly incorporates geometry conditions to enhance the model’s awareness of scene structure and depth continuity. Specifically, we utilize depth and surface normal videos as geometric conditions. These two types of geometric cues are complementary — depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations. When combined, they offer a richer and more holistic description of scene geometry (see Sec. 3.2 for details on data acquisition). Since both the depth and normal videos are spatially aligned with the RGB views, we propose GuideNet, which utilizes a stack of convolutional layers that jointly downsample and encode the depth and normal sequences. The resulting geometry-aware representation is then concatenated with the noise latents along the channel dimension, allowing the generation process to be guided by accurate 3D structural cues.
@@
Appearance Conditions. To enable fine-grained control over the generated textures, RoboTransfer introduces texture conditions from two complementary perspectives: background appearance and object appearance. Specifically, we use a background reference image and a set of object reference images as conditions signals to guide the generation process toward faithful reproduction of scene textures and object-level details. A key challenge in appearance control is ensuring compatibility with the geometry conditions. Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the structural consistency and the visual fidelity of generated results. To address this, we carefully curate the appearance reference images, resulting in a background reference image $C_{b}$ and object reference images $C_{o}$ that do not contradict the geometric priors (see Sec. 3.2 for details). For background appearance, we employ a VAE encoder to compress $C_{b}$ into a latent representation that is spatially aligned with the generation latents. This encoded background appearance is then concatenated with the latent inputs to control the global texture and background style of the output video. In contrast, object appearance presents additional challenges due to the variable number and spatial distribution of objects in the scene. Therefore, object images are treated as unstructured conditions. We encode each object’s appearance using CLIP [47], which produces a global embedding used to guide the generation via cross attention. This design allows flexible and scalable conditioning on diverse object appearances while preserving compatibility with the structured geometry inputs.
[…]
Figure 3: RoboTransfer data processing pipeline. The data processing pipeline consists of two main components: Geometry conditions (left) are derived by using Mono-Normal and Video Depth Anything, which are scale-aligned with sensor depth to ensure consistency. Appearance conditions (right) are obtained by sampling keyframes from the video. GPT-4 is used to generate object descriptions, which are then processed by Grounding-SAM to create per-object masks. Additionally, background inpainting is used to generate complete reference backgrounds.
3.3 Dataset Construction from Real World Data
Data construction is central to our framework. The overall pipeline is illustrated in Figure 3. Unlike conventional text-to-video or image-to-video tasks, our approach requires both globally consistent geometry and multiple part-specific reference images (e.g., background and manipulable objects).
Our pipeline begins with the collection of a large-scale robotic demonstration dataset. Each demonstration captures synchronized RGB-D streams from both a hand-eye and an overhead RGB-D camera. To cover a wide range of operating conditions, we vary backgrounds (e.g., textured tabletops, cluttered surfaces) and manipulable objects (shapes, sizes, materials).
To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals. Since raw depth maps from RGB-D sensors are often noisy, we employ the state-of-the-art video depth estimator, Video Depth Anything[9], to produce temporally consistent depth maps. We then align the scale of the estimated depth map with the RGB-D sensor’s depth using robust least-squares fitting to ensure global spatial accuracy. Additional details are provided in the supplementary material. For surface normals, we compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS[24], which jointly serves as reliable geometric conditions.
Next, we sample a subset of RGB keyframes as visual references. These frames are processed by GPT-4 to generate structured scene descriptions, including object-level descriptions for individual entities. These descriptions guide the Grounding-SAM module[50], which detects and segments per-object masks. To obtain clean background references, we inpaint the segmented objects from the original images, generating plausible object-free scenes (e.g., an empty tabletop). For multimodal object-level conditions, the segmented object masks are further processed with the CLIP model[47] to obtain semantic embeddings. These conditions allow the diffusion model to generate photorealistic, semantically grounded video augmentations that respect both object identity and spatial layout.
4 Experiments
4.1 Synthesis Quality
Implementation Details. To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions, covering diverse tasks such as grasping and placing. The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training. Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation. All videos are annotated with conditions as described in Sec. 3.2, to facilitate both training and evaluation. RoboTransfer is fine-tuned from the pretrained video diffusion model [5]. During training, videos from each camera view are resized to a resolution of $640\times 384$. We adopt the AdamW optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps. During inference, we use the EDM scheduler [29] to perform 30 denoising steps and apply classifier-free guidance.
Table 1: Experiment comparison on different geometry conditions of RoboTransfer.
@@ 4.1 Synthesis Quality
RoboTransfer [Scaled D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
RoboTransfer [Scaled D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
RoboTransfer [D.S.] middle 0.182 0.074 0.031 3.51 1.58 153.76
RoboTransfer [D.P.] middle 0.132 0.053 0.015 3.05 1.43 95.89
RoboTransfer [Scaled D.P.] middle 0.134 0.054 0.015 2.93 1.40 103.32
RoboTransfer [Scaled D.P. + N.] middle 0.133 0.054 0.015 2.86 1.39 101.17
@@
RoboTransfer [Scaled D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
RoboTransfer [Scaled D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
Geometry Consistency Analysis. We first validate the geometric consistency in RoboTransfer’s generative capability. As shown in Table 1, directly using raw depth sensor data (D.S.) as a conditional input introduces considerable noise, which adversely affects the generation quality. This degradation is evident across all three views in terms of depth accuracy, surface normals, and multi-view consistency. In contrast, employing model-predicted [9] depth (D.P.) results in smoother depth maps, which serve as a more stable supervisory signal for training the generative model. Consequently, the generated videos exhibit significantly better geometric consistency. Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improves the RMSE and Mean Err. of the middle view by 27.4% and 14.2%, and increases Pix.Mat. of the middle and left/right views by 4.7% and 37.6%, respectively. Further improvements are observed where we incorporate scaled predicted depths (Scaled D.P.) to enforce a consistent depth scale across multiple views. This multi-view scale alignment further enhances geometric consistency, leading to relative improvements of 22.4% and 16.8% in Pix.Mat. of left and right views. Finally, combining normal maps with depth as joint conditions (Scaled D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
Appearance Consistency Analysis.
@@
middle 0.712 0.847 0.137 1.46 108.65
middle 0.719 0.845 0.135 1.41 105.66
middle 0.712 0.855 0.136 1.42 105.54
middle 0.720 0.858 0.133 1.39 101.17
@@
We then evaluate RoboTransfer’s capability in controlling visual appearance. As shown in Table 2 (row-1 and row-2 of all three camera views), we observe that directly conditioning on background images without inpainting can corrupt the geometric cues. In contrast, inpainting the background before conditioning not only preserves the geometric structure but also improves appearance consistency, leading to $\sim 1\%$ relative improvements in RMSE, median error, and background similarity. Comparing row-1 and row-3 of the middle camera view, we examine object-level appearance control. In row-1, the entire image with all the objects is encoded using a global CLIP feature. In contrast, row-3 applies a finer-grained approach by masking and encoding each object individually, followed by feature concatenation. This object-wise conditions leads to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances. Finally, row-4 of all three camera views demonstrate that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control. This setup yields the highest CLIP similarity scores for both regions, confirming the effectiveness of jointly modeling structured and unstructured visual components.
Visualizations. We demonstrate that RoboTransfer enables controllable manipulation of both foreground and background appearance while preserving geometric and multi-view consistency across camera views. As shown in Figure 4, RoboTransfer successfully modifies the background appearance while retaining the foreground object’s texture and scene geometry. Figure 5 illustrates that the model can alter the foreground appearance without affecting the background or spatial layout. Finally, Figure 6 demonstrates the model’s ability to jointly edit both foreground and background elements. All visualizations confirm that RoboTransfer maintains geometric and multi-view consistency throughout the generated scenes.

### v2

# RoboTransfer: Controllable Geometry-Consistent Video Diffusion for Manipulation Policy Transfer

@@ Abstract
The goal of general-purpose robotics is to create agents that can seamlessly adapt to and operate in diverse, unstructured human environments. Imitation learning has become a key paradigm for robotic manipulation, yet collecting large-scale and diverse demonstrations is prohibitively expensive. Simulators provide a cost-effective alternative, but the sim-to-real gap remains a major obstacle to scalability. We present RoboTransfer, a diffusion-based video generation framework for synthesizing robotic data. By leveraging cross-view feature interactions and globally consistent 3D geometry, RoboTransfer ensures multi-view geometric consistency while enabling fine-grained control over scene elements, such as background editing and object replacement. Extensive experiments demonstrate that RoboTransfer produces videos with superior geometric consistency and visual fidelity. Furthermore, policies trained on this synthetic data exhibit enhanced generalization to novel, unseen scenarios. Project page: https://horizonrobotics.github.io/robot_lab/robotransfer.
[…]
Figure 1: Overview of RoboTransfer. While real-world data collection is prohibitively expensive and simulated data often lacks visual fidelity, RoboTransfer synthesizes high-quality data with strict multi-view consistency. Our experiments demonstrate that this synthetic data significantly enhances policy performance on real robots.
11footnotetext: Equal contribution.22footnotetext: Corresponding author: nemo.liu@horizon.auto, zhizhong.su@horizon.auto
Table 1: Comparison of different methods for robotic data generation. RoboTransfer is the only one that provides temporal and multi-view consistency while offering fine-grained control.
Method Model Type Generation Consistency Control Level
Temporal Multi-View Background Object Environment
ROSIE [54] Image Diffusion
RoboEngine [55] Image Diffusion
Cosmos-Transfer1 [2] Video Diffusion
RoboTransfer (Ours) Video Diffusion
@@ 1 Introduction
Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [60]. However, collecting large-scale real-world robot demonstrations is prohibitively expensive [6, 7]. Simulated environments offer a cost-effective alternative [51, 37], but the scarcity of assets and sim-to-real gap[43, 36] make it extremely challenging to scale.
@@
Despite rapid advancements, two key challenges persist. First, robotic systems often rely on single-view observations, limiting their perception. Multi-view observation is commonly used to address this, but generating consistent multi-view results remains a significant challenge for video generative models. Second, robot manipulation tasks are complex and interactive, and precisely controlling these tasks via textual input alone, as seen in text-to-video frameworks [5, 63], remains a substantial challenge.
To address these challenges, we introduce RoboTransfer, a geometry-consistent video diffusion framework tailored for robotic visual policy transfer (Figure 1). As summarized in Table 1, RoboTransfer is, to the best of our knowledge, the first framework to guarantee multi-view generation consistency in robotic data synthesis, while also offering fine-grained and disentangled control over both background and object attributes.
The main contributions are as follows:
@@
    We propose RoboTransfer, a video data generation framework for robotic manipulation that ensures multi-view consistency while enabling fine-grained, disentangled control.
@@
    We introduce a novel data construction pipeline that automatically decomposes real-world robot demonstrations into the geometric and appearance conditions.
@@
    We demonstrate that RoboTransfer generates multi-view videos with substantially improved generation consistency, and that policies trained on this data generalize more effectively to novel, unseen environments.
@@ 2.1 Visual Generalizable Imitation Learning
Imitation Learning (IL) has become a cornerstone for visuomotor control in robotic manipulation, with deep networks mapping raw visual inputs to motor commands based on demonstrations [10, 26, 60, 7]. However, collecting large-scale task-specific data is prohibitively expensive, leading to the use of auxiliary sources such as human demonstration videos [17, 4] and operational logs from other platforms [39, 25, 8]. These sources introduce domain-specific biases and distributional shifts, which can degrade IL policy performance when transferred to new environments. While simulators can generate large quantities of labeled frames [45, 51, 16, 37], discrepancies in physics modeling, rendering fidelity, and scene composition hinder sim-to-real transfer. Domain randomization attempts to close this gap [30, 18, 29, 1], but it typically only applies color perturbations across the entire image, without capturing localized or structurally meaningful variations. In this work, we leverage a generative model to synthesize photorealistic multi-view videos for manipulations, significantly enhancing policy robustness and enabling seamless transfer to novel environments.
[…]
Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints. It represents geometry with metric depth and normal maps, and encodes appearance using reference backgrounds and object images for detailed control over appearance.
@@ 2.2 Generation Models in Embodied AI
With the rapid advancement of generative models  [28, 47, 5, 2, 53, 63, 13, 56, 31, 19], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [65]. In particular, domains like autonomous driving have seen increasing adoption of generative techniques, as demonstrated by [49, 59, 22, 50, 14, 15, 57, 38, 58]. Similarly, in robotic scenarios, where real-world data collection is often prohibitively expensive and time-consuming, generative data has proven to be highly beneficial. Models such as UniSim [52], UniPi [11], and RoboDreamer [64] synthesize future robot behaviors via text prompts, which are then translated into actionable commands using inverse dynamics models. However, purely video-based generation methods often suffer from spatial and temporal inconsistencies, leading to degraded performance in downstream action prediction and planning. To address this, TesserAct [62] proposes a multi-modal generation pipeline that simultaneously synthesizes RGB, depth, and normal videos, enabling the reconstruction of coherent 4D scenes. This framework ensures both spatial and temporal consistency in robotic environments and supports policy learning that significantly surpasses prior video-only world models. Moreover, data generalization remains a critical challenge in robot learning. Techniques like ReBot [12] and RoboEngine [55] adopt background inpainting to diversify environmental textures, while Cosmos-Transfer [2] utilizes video-to-video translation to enrich overall scene appearance. Despite their success, these approaches lack fine-grained control over foreground and background textures, often resulting in synthetic data distributions misaligned with real-world tasks, especially when training policy models. Additionally, their inability to ensure multi-view consistency limits their scalability and effectiveness in complex, robotic settings.
@@ 3.1 Video Diffusion Model Preliminaries
Diffusion-based methods for controllable video generation model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding). Popular approaches in this paradigm include Flow Matching [32], DDPM [21], and EDM [24]. In the case of EDM, the learning objective is formulated as:
$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$
@@
$\vskip-2.0pt\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
@@
$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).\vskip-2.0pt$ (4)
@@ 3.2 RoboTransfer Framework
Figure 3: RoboTransfer data construction pipeline generates image pairs with Geometry and Appearance conditions. Geometry conditions (left) with metric depth and normals from demonstration videos and appearance conditions (right) from keyframes. A VLM-based descriptor generates object descriptions, which are processed by Grounding-SAM to create per-object masks.
@@
The overall framework of RoboTransfer is illustrated in Figure 2. To ensure multi-view consistency during generation, we perform multi-view consistent modeling, enabling the generation process to reason jointly over information from different viewpoints. On the conditions side, RoboTransfer incorporates fine-grained control by encoding both geometric and appearance information. Specifically, we represent geometry using metric depth maps and surface normal maps, capturing the underlying 3D structure of the scene. Meanwhile, the appearance is encoded using reference background images and object-specific images, providing detailed control over texture, color, and contextual appearance. In the following sections, we first introduce the multi-view consistent modeling and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
@@
Multi-view Consistent Modeling. Robotic manipulation often relies on multi-view camera setups to capture a scene in parallel. To generate multi-view consistent videos for downstream tasks, RoboTransfer leverages the multi-view in-context learning capabilities [59, 23] inherently present in pretrained diffusion models. Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a Variational Autoencoder(VAE) encoder $\mathcal{E}$:
$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),\vskip-2.0pt$ (5)
@@
where $V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence. This modeling approach leverages the spatial reasoning capabilities of existing video diffusion backbones, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic manipulation settings.
@@
Geometry Conditions Injection. Recent video generation models primarily learn spatiotemporal coherence through data-driven approaches such as temporal attention [5, 63] or full attention [47, 28, 2]. While effective at capturing pixel-level dynamics, these methods lack an explicit understanding of underlying 3D geometry, limiting their applicability in robotic environments. To address this limitation, RoboTransfer explicitly incorporates geometry conditions to enhance the model’s awareness of scene structure and depth continuity. Specifically, we utilize depth and surface normal videos as geometric conditions. These two types of geometric cues are complementary; depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations. When combined, they offer a richer and more holistic description of scene geometry (see Sec. 3.3 for details on data acquisition). Since the depth and surface normal sequences are spatially aligned with their corresponding RGB views, we leverage the VAE encoder composed of stacked convolutional layers that jointly downsample and encode geometric cues from depth and normal videos. The resulting geometry-aware representation is concatenated with the noise latents along the channel dimension, enabling diffusion-based video generation to be precisely guided by consistent and physically plausible 3D cues throughout the generation process.
@@
Appearance Conditions Injection. To enable fine-grained control over the generated textures, RoboTransfer introduces texture conditions from two complementary perspectives: background appearance and object appearance. Specifically, we use a background reference image and a set of object reference images as condition signals to guide the generation process toward faithful reproduction of scene textures and object-level details. A key challenge in appearance control is ensuring compatibility with the geometry conditions. Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the geometry consistency and the visual fidelity of generated results. To address this, we carefully curate the appearance reference images, resulting in a background reference image $C_{b}$ and object reference images $C_{o}$ that do not contradict the geometric priors (see Sec. 3.3 for details). For background appearance, we employ a VAE encoder to compress $C_{b}$ into a latent representation that is spatially aligned with the generation latents. This encoded background appearance is then concatenated with the latent inputs to control the global texture and background style of the output video. In contrast, object appearance presents additional challenges due to the variable number and spatial distribution of objects in the scene. Therefore, object images are treated as unstructured conditions. We encode each object’s appearance using CLIP [40], which produces a global embedding used to guide the generation via cross attention. This design allows flexible and scalable conditioning on diverse object appearances while preserving compatibility with the structured geometry inputs.
Table 2: Experiment comparison on different geometry conditions of RoboTransfer: It shows that combining metric depth and normal maps yields the best consistency across all views and metrics.
@@
RoboTransfer [Metric D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
RoboTransfer [Metric D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
RoboTransfer [D.S.] head 0.182 0.074 0.031 3.51 1.58 153.76
RoboTransfer [D.P.] head 0.132 0.053 0.015 3.05 1.43 95.89
RoboTransfer [Metric D.P.] head 0.134 0.054 0.015 2.93 1.40 103.32
RoboTransfer [Metric D.P. + N.] head 0.133 0.054 0.015 2.86 1.39 101.17
@@
RoboTransfer [Metric D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
RoboTransfer [Metric D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
@@
head 0.712 0.847 0.137 1.46 108.65
head 0.719 0.845 0.135 1.41 105.66
head 0.712 0.855 0.136 1.42 105.54
head 0.720 0.858 0.133 1.39 101.17
@@
3.3 Dataset Construction from Real Robotic Demonstration Data
Data construction is central to our framework, where real-world recording data is decomposed into high-quality triplets for training Robotransfer, consisting of Geometry Conditions, Appearance Conditions, and ground-truth images. The overall pipeline is illustrated in Figure 3.
Geometry Conditions Construction Starting from robotic demonstration data with two wrist cameras and one head camera. To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals. Since some robots only have RGB sensors and raw depth maps from RGB-D sensors are often noisy, we use a state-of-the-art depth estimator [9, 48] to produce consistent depth maps. For unmetric depth outputs, we align the estimated scale with the RGB-D sensor using robust least-squares fitting to ensure global spatial accuracy. For surface normals, we compute per-frame estimates using state-of-the-art monocular normal estimators [20, 48], offering conditions for detailed geometry.
Appearance Conditions Construction. To enable controllable data generation with the diffusion model, we sample multiview RGB keyframes as appearance conditions. A VLM-based descriptor generates scene and object descriptions, which guide the Grounding-SAM module [42] to detect and segment per-object masks. To obtain clean background references, the segmented objects are inpainted in the original images, producing plausible object-free scenes (e.g., an empty tabletop). For object-level conditions, the segmented masks are further processed with the CLIP model [40] to extract semantic embeddings.
4 Experiments for Synthesis Quality